what is the smallest possible slope of y = x^3-3x^2+5x
Answer by Alan3354(69443)

y = x^3-3x^2+5x
y' = 3x^2-6x+5
y" = 6x - 6 = 0
x = 1
y'(1) = 3-6+5 = 2
2 is the minimum slope, at (1,3) and is a point of inflection